OpenStudy (anonymous):

Find the number of terms in the sequence. -10,-6,-2,...........38

OpenStudy (anonymous):

here's hint If you have a number sequence then the numbers in the sequence are called the terms. So if you are asked for the term to term rule then you are looking for a rule that takes you from the previous term in the sequence to the next term in the sequence. For example, if you have a number sequence that goes, 2,5,8,11,14 then the term to term rule is add on 3 (since its increasing by 3 each time). This can be written down more formally as: xn+1 = xn + 3 xn+1 is the next term in the sequence. xn is the term that you are currently on Or if you have another sequence that goes 13,11,9,7,5 then the tem to term rule is take away 3 (since the sequence is decreasing by 3 each time). Again this can be written using as algebra as: xn+1 = xn - 2 Finding the position to term rule for a number sequence is a little harder (also known as the nth term). The first term in the sequence is in position 1, the second term is in position 2, the third term is in position 3, the fourth term is in position 4, and the fifth term is in position 5. So the position to term rule is the rule that you takes from the position values to the term in the sequence. For example, a sequence goes 5,9,13,17,21... then the position to term rule is to multiply the position number by 4 and add on 1. Since: 1 × 4 + 1 = 5 (since 5 is the 1st number in the sequence). 2 × 4 + 1= 9 (since 9 is the 2nd number in the sequence). 3 × 4 + 1 = 13 (since 13 is the 3rd number in the sequence). 4 × 4 + 1 = 17 (since 17 is the 4th number in the sequence). 5 × 4 + 1 = 21 (since 21 is the 5th number in the sequence). The rule can be written using algebra as 4n + 1 (where n is the position number). Or if your sequence goes 1,8,15,22,29... then the position to term rule is to multiply by 7 and take off 6. Since: 1 × 7 - 6 = 1 (since 1 is the 1st number in the sequence). 2 × 7 - 6= 8 (since 8 is the 2nd number in the sequence). 3 × 7 - 6 = 15 (since 15 is the 3rd number in the sequence). 4 × 7 - 6 = 22 (since 22 is the 4th number in the sequence). 5 × 7 - 6 = 29 (since 29 is the 5th number in the sequence). The rule can be written using algebra as 7n – 6

OpenStudy (lgbasallote):

d = 4 A1 = -10 An = 38 An = A1 + (n-1)d plug in the values...you're looking for n

OpenStudy (anonymous):

@roh, i think you have one of those transcriber thingys and your compter types what you say. :)

OpenStudy (lgbasallote):

i knew he wasn't typing fast! *gasps*